# Frostman random variables, entropy inequalities, and applications

### v4

@@ 7 Examples
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculations reveal that
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll M(r_{1})M(r_{2})\cdot 2^{4m}$
where
@@
In all cases, we have
@@
for any $n\in\mathbb{N}$ and any random variable $X$ with bounded range, where $Y$ is i.i.d. with $X$. In this example, we show that the above statement may fail when $k,l$ are $\mathbb{Q}$-independent. In particular, we consider the situation that $(k,l)=(1,\sqrt{2})$.
@@
One can deduce by Lemma 2.19 that $H_{n}(X+\sqrt{2}Y)=H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})+O(1).$ So, it is sufficient to prove that $H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=2m$.
Note that both $X^{\prime}$ and $Y^{\prime}$ have discrete uniform distribution on $A^{\prime}$. Denote by $p_{j}=P\big(D_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=j\big)$ for $0\leq j<(1+\sqrt{2})2^{n}$. By Claim 1, we see that either $p_{j}=2^{-2m}$ or $p_{j}=0$. Therefore,
@@
Appendix A Appendix - Existence of Certain Random Variable
@@
$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}).$
Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, if and only if $h(Z(\omega))=h(Z(\nu))$. The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
@@ Acknowledgements
We would like to thank Tuomas Orponen for his comments on the first version of this paper and for telling us about the references [1, 22].

### v5

@@ 8 Examples
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll$
$\qquad\begin{cases}r_{1}\cdot r_{2}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-4m},\\ r_{1}\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m}<r_{2}\leq 2^{-m},\\ r_{1}\cdot 2^{-4m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m},\,2^{-m}<r_{2}\leq 1,\\ r_{1}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m},\,r_{2}>1,\end{cases}$
$\qquad\begin{cases}2^{-4m}\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq r_{2}\leq 2^{-3m},\\ 2^{-4m}\cdot 2^{-4m}\frac{r_{2}}{2^{-3m}}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m}<r_{2}\leq 2^{-2m},\\ 2^{-4m}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,2^{-2m}<r_{2}\leq 2^{-m},\\ 2^{-4m}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,2^{-m}<r_{2}\leq 1,\\ 2^{-4m}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,r_{2}>1,\end{cases}$
$\qquad\begin{cases}2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}\frac{r_{2}}{2^{-3m}}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq r_{2}\leq 2^{-2m},\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m}<r_{2}\leq 2^{-m},\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m},\,2^{-m}<r_{2}\leq 1,\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m},\,r_{2}>1,\end{cases}$
$\qquad\begin{cases}2^{-4m}2^{m}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq r_{2}\leq 2^{-m},\\ 2^{-4m}2^{m}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m}<r_{2}\leq 1,\\ 2^{-4m}2^{m}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}>1,\end{cases}$
@@
The rest of the cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
@@
for any $n\in\mathbb{N}$ and any random variable $X$ with bounded range, where $Y$ is i.i.d. with $X$. In this example, we show that the above statement may fail when $k,l$ are $\mathbb{Q}$-independent real numbers. In particular, we consider the situation that $(k,l)=(1,\sqrt{2})$.
@@
One can deduce by Lemma 2.19 that $H_{n}(X+\sqrt{2}Y)=H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})+O(1).$ So, it is sufficient to prove that $H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=2m$. Note that both $X^{\prime}$ and $Y^{\prime}$ have discrete uniform distribution on $A^{\prime}$. Denote by $p_{j}=P\big(D_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=j\big)$ for $0\leq j<(1+\sqrt{2})2^{n}$. By Claim 1, we see that either $p_{j}=2^{-2m}$ or $p_{j}=0$. Therefore,
@@
Appendix A Appendix - Existence of Certain Random Variables
@@
$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}),$
which is a disjoint union. Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, which is also equivalent to $h(Z(\omega))=h(Z(\nu))$. The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartesian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
@@ Acknowledgements
We are grateful to Tuomas Orponen for a series of insightful questions and valuable comments on the first version of this paper, which led us to substantially strengthen the results and greatly improve the exposition. In particular, he pointed out that the discretized sum-product problem for $\phi(x,y)=x^{2}+y^{2}$ and a complete graph $G$ could be approached using the techniques from [6], [30], [26], [27].